OpenStudy (anonymous):

Divide x to the 7 ninths power over x to the 1 fourth power

OpenStudy (anonymous):



OpenStudy (anonymous):

When you're dividing by an exponent, you subtract the exponent on the bottom from the exponent on the top. To do this, you must first convert both fractions into ones with equal denominators. It should be their least common multiple (36).

OpenStudy (anonymous):

\[7 \times 4 \over 9 \times 4\] \[1 \times 9 \over 4 \times 9\]

OpenStudy (anonymous):

\[9 \over 36\]\[28 \over 36\]